Please help me. We are working on probability and I am not so good with it. I am not sure it the answer would be 1/23--that just seems too easy. If you can please help me, I would be thankful. Here is the problem.
A bag contains 9 green marbles, 8 red marbles, and 7 blue marbles. Without looking, a marble is drawn from the bag. What is the probability that a red or a blue marble will be drawn?

Answer by checkley75(3666)

green+red+blue=total.
9+8+7=24 total marbles.
(8+7)/24=15/24=5/8=.625 or 62.5% chance that a red or a blue marble will be drawn.